OpenStudy (anonymous):

suppose two lines l1 ad l2 are parallel to vectors i+j and i-2j +k then find angle between l1 and l2?

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@hari5719

OpenStudy (anonymous):

@RANE

OpenStudy (michele_laino):

Please note that, using an orthogonal coordinate system, we can write: \[i+j=(1,1,0),\quad i-2j+k=(1,-2,1)\] so please use the canonical scalar product, in order to find the cosine function of the angle enclosed between your lines

OpenStudy (anonymous):

understood nothing

OpenStudy (michele_laino):

If we have two vectors, namely a and b, we can know the angle theta enclosed by them using the subsequent formula: \[\cos \theta =\frac{ a \cdot b }{ |a||b| }\] where |a|, and |b| are the lengths of vectors a and b respectively

OpenStudy (anonymous):

ok.thanks

OpenStudy (michele_laino):

thanks